Write down the condition under which a large molecular cloud collapses to give rise to new stars. Calculate the time of free fall of an HI cloud of density (10)6 particles/(m)3.
1
Expert's answer
2019-04-03T09:11:59-0400

Write down the condition under which a large molecular cloud collapses to give rise to new stars.

Solution:

For collapse of a cloud withradius R to occur, need either


"M_{cloud}>M_{Jeans}=\frac{3kTR}{2G \mu m_H}"




Calculate the time of free fall of an HI cloud of density (10)6 particles/(m)3.

Solution:

The acceleration g felt by a test particle for a spherically symmetric distribution of mass M and radius r.

From Newton's second law, the equation of motion for a test particle at the edge of the cloud is then


"m \times g=G \times m \frac{M}{r^2} (1)"

The mass M is equal to:


"M= \frac{4 \pi}{3}\times r^3 \times ρ (2)"

where ρ is the cloud density


We put (2) in (1):


"g=\frac{4 \pi}{3}G \times r \times \rho (3)"

If it starts initially at rest, then (if acceleration is constant) it will reach the center when


"\frac{ g \times t^2}{2}=r (4)"

We put (3) in (4) and solve for t:


"t=\sqrt \frac {3}{2\times \pi \times G \times \rho} (5)"

The cloud density ρ is equal to:


"\rho=n \times m_H (6)"

where n=106 particles/m3 ; mH = 1.67 × 10-27 kg

Using (6) we calculate the value of cloud density ρ: ρ= 1.67 × 10-21 kg/m3

We put the value of cloud density ρ in (5) and get:

t=2× 1015 seconds

One year contains 3× 107 seconds

In this case t=67.000.000 years


Answer:

67.000.000 years